If James needs $15,000 for his new car. How much does he have to invest now in a 36-month CD that has a rate of 3.75% compounded quarterly?
Answer by stanbon(48558)

If James needs $15,000 for his new car. How much does he have to invest now in a 36-month CD that has a rate of 3.75% compounded quarterly?
--------------
15000 = P(1 + (0.0375/4))^(4*3)
15000 = P(1.009375)^(12)
P = 15000/1,118486
P = $13,410.99
=================
Probably the answer is $13,411.00
=================
Cheers,
Stan H.